# One-step Noisy Label Mitigation

### v1

@@ 4.2 Comparisons with State of The Arts
Results on Flickr30K.
To further demonstrate the generalization ability of OSA, we evaluate on the Flickr30K dataset and compare with several anti-noise methods, including NCR [1], DECL [9], BiCro [7], and NPC [2]. The results are presented in Table. 3 of Appendix. It is evident that OSA consistently outperforms all models on the R@1 metric. Notably, compared with the baseline CLIP, training with OSA at a 60% noise ratio achieves 20.9% R@1 improvement for i2t and a 22.3% R@1 improvement in t2i, further indicating the effectiveness of OSA on noise mitigation. Additionally, OSA demonstrates similar noise robustness on the Flickr30K dataset as observed on MSCOCO, with only 1.4% R@1 drop on i2t and 1.2% R@1 drop on t2i ranging from 0% noise to 60% noise, while all of the other anti-noise approaches hardly resist the detriment from high-ratio noise. All of these results demonstrate the effectiveness and robustness of OSA on anti-noise tasks.
Results on CC120K.
To further verify the reliability of OSA in real scenarios, we conduct evaluations on a large-scale real-world noisy dataset, CC120K, with 3%-20% noise ratio. The results shown in Table. 5 indicate that OSA outperforms the current state-of-the-art method NPC, even in larger-scale real-world domains. This demonstrates the feasibility and generality of OSA even in practical training scenarios.
Table 4: Comparison on real-world noisy dataset CC120K.
i2t t2i
Method R@1 R@5 R@10 R@1 R@5 R@10
NPC 71.1 92.0 96.2 73.0 90.5 94.8
CLIP 68.8 87.0 92.9 67.8 86.4 90.9
+OSA 73.1 92.2 95.7 73.9 91.2 94.7
Table 5: Results of other image-based tasks.
Method Image Classification Image Retrieval
Aircraft Bird Car Prec. mAP
Acc Acc Acc
Baseline 65.44 62.29 75.90 71.69 18.16
+OSA 73.18 70.50 80.19 78.45 24.99
Results on Other Downstream Tasks.
To validate the transferability of OSA across different tasks, we evaluate it on two additional tasks: image classification and image retrieval. The results are presented in Table. 5. The baseline method for both tasks leverages contrastive learning. In the image classification task, OSA outperforms the baseline by 7.74%, 8.21%, and 4.28% on the Aircraft, Bird, and Car subsets, respectively. In the image retrieval task, OSA improves performance by 6.76% in precision and 6.83% in mAP. These improvements demonstrate the strong task transferability and generality of OSA.
4.3 Target Model-Agnostic Analysis
OSA is an architecture-agnostic paradigm that can be easily adapted to various models. To verify its model-agnostic property, we evaluate it across models with different architectures. Subsequently, we apply it to other anti-noise models to demonstrate its generalization capability in noise mitigation.
Architecture-agnostic Analysis.
The effectiveness of OSA on Vision Transformer (ViT) has been proven in Section. 4.2. We further explore the generality of OSA on target models with other architectures. Specifically, we deploy OSA above the VSE++ [21] model with two different architecture types: ResNet-152 [22] and VGG-19 [23]. These two architectures have revealed significant sensitivity and vulnerability to noise [1]. In this experiment, all estimator models employ zero-shot CLIP and we utilize the original VSE++ as our baseline. The results in Table. 6 indicate a significant performance degradation emerged for the baseline methods in noisy setting, while a stable performance is achieved after employing OSA. The stable performance on these two noise-vulnerable architectures fully demonstrates that OSA possesses the architecture-agnostic property.
Table 6: The results of the target model with different architectures on noisy MSCOCO.
Noise ratio Method Architecture MS-COCO 1K MS-COCO 5K
@@ 4.3 Target Model-Agnostic Analysis
0% Baseline ResNet-152 58.9 86.9 93.8 44.2 77.9 88.3 34.9 64.3 76.1 23.3 50.9 64.2
  +OSA 58.9 86.2 93.7 44.3 77.9 87.9 35.0 64.1 76.0 23.5 50.8 63.9
Baseline VGG-19 49.6 79.4 89.1 38.0 72.9 84.7 26.9 54.2 66.8 18.7 43.8 56.8
  +OSA 50.1 80.0 89.3 38.3 73.0 84.6 26.6 54.4 67.4 18.8 43.9 57.3
20% Baseline ResNet-152 45.8 70.3 83.7 36.1 68.4 79.7 26.0 48.4 58.3 18.3 42.0 54.0
  +OSA 58.1 86.1 93.2 43.4 76.8 87.2 33.7 62.6 74.5 22.5 49.7 62.8
Baseline VGG-19 33.2 67.1 81.5 25.9 58.0 71.4 13.7 35.0 49.2 10.7 29.9 41.9
  +OSA 49.3 79.1 88.6 37.2 71.9 83.8 25.2 53.3 65.3 17.9 42.6 55.9
50% Baseline ResNet-152 28.4 61.2 75.2 5.2 14.0 19.5 11.0 31.0 43.6 1.6 6.0 9.2
  +OSA 55.0 84.0 92.0 40.7 74.7 85.6 30.8 60.2 72.3 20.9 46.6 60.0
Baseline VGG-19 2.5 9.8 16.2 0.1 0.5 1.0 0.5 2.5 4.4 0.0 0.1 0.2
  +OSA 47.1 77.7 87.6 35.7 70.3 82.8 24.0 51.5 64.0 16.9 40.8 54.2
@@
i2t t2i
Noise Ratio Method R@1 R@5 R@10 R@1 R@5 R@10
@@
Adaptability to Other Anti-Noise Models.
Theoretically, OSA can be adapted to any target model, providing noise resistance. However, can OSA further enhance the robustness of models specifically designed for noise mitigation? To investigate this, we applied OSA to the current state-of-the-art model, NPC [2]. As shown in Table. 7 of Appendix, even for noise-mitigating models, OSA consistently improves training robustness. This finding further demonstrates the broad adaptability of OSA across different model types.
4.4 Estimator Model Analysis.
The estimator model is the basis of OSA’s anti-noise capability. In this section, we explore the impact of different estimator models on noise mitigation, and examine the impact of domain adaptation in noise mitigation. In Table. 8 of Appendix, we investigate four types of estimators: “None” refers to training CLIP directly without using OSA. “CLIP (w/o DA)” and “ALIGN (w/o DA)” represent using CLIP and ALIGN without domain adaptation as estimators, respectively, i.e., zero-shot CLIP and ALIGN. “CLIP (w DA)” indicates the CLIP with domain adaptation. The target models are all CLIP. We can observe that both of CLIP and ALIGN as estimators significantly enhance the target model performance stability when learning with noise, indicating that the choice of estimator is very flexible. Both CLIP and ALIGN demonstrate exceptional performance when served as estimators. The other phenomenon is that the zero-shot CLIP model shows comparable performance to the domain-adapted CLIP with a even better performance at lower noise ratios. This indicates that zero-shot CLIP, as an estimator, already performs exceptionally well in noise mitigation. The domain adaptation is unnecessary. This further enhances the deployment convenience of OSA.
Table 8: Ablation study of estimator type on noisy MS-COCO.
@@
CLIP (w DA) 80.4 96.2 98.6 67.8 91.6 96.4 64.0 85.5 91.9 47.9 74.6 83.8
4.5 Noise Assessment Accuracy
Noise Detection Accuracy Analysis.
To figure out how accurate OSA is in recognizing noise, we evaluate the accuracy and recall on CLIP without Domain-Adaptation (w/o DA) and CLIP with Domain-Adaptation (w DA) on noisy MSCOCO. We utilize zero as the threshold to roughly divide pairs into noise and clean sets, respectively. Concretely, we classify scores less than or equal to 0 as noise, and scores greater than 0 as clean. The Accuracy means the proportion of the clean pairs correctly classified into the clean set, while the Recall indicates the noisy pairs correctly classified into the noisy set. The results presented in Table. 11 indicates the powerful noise recognizing capability of OSA. The remarkable performance on CLIP (w/o DA) fully demonstrates the generality of OSA. Another notable phenomenon is that all recall scores converge towards 100, indicating that OSA achieves greater accuracy in noise detection. This suggests that OSA can almost entirely eliminate the impact of noise on training.
@@ 4.5 Noise Assessment Accuracy
Noise Re-weighting Accuracy Comparison.
Some anti-noise methods, like NPC, also employ loss re-weighting for optimization. To assess whether our method assigns relatively smaller weights to noise than these methods, we first analyze the weights generated by NPC and OSA. Due to differences in weight scales across methods, a direct comparison is unfair. Therefore, to unify the scale, we adopt a ranking-based approach, sorting weights in descending order and calculating the Mean Noise Rank. This metric evaluates whether smaller weights are consistently assigned to noisy samples relative to clean ones. Our experiments use 2,000 randomly selected samples from the MSCOCO dataset under two noise conditions: 20% noise (370 noisy samples) and 50% noise (953 noisy samples). The theoretical optimal Mean Noise Ranks, where all noisy weights are ranked last, are 1815.5 and 1524.0, respectively. Results presented in Table. 9 of Appendix show that OSA achieves a higher Mean Noise Rank compared to NPC, demonstrating greater accuracy in re-weighting. Moreover, OSA’s rankings are nearly optimal (20% noise: 1809.1 for OSA versus 1815.5 optimal; 50% noise: 1520.7 for OSA versus 1524.0 optimal). This near-perfect alignment indicates that OSA effectively places almost all noisy samples behind the clean ones.
Table 10: ACC and recall of noise detection.
Estimator Type Noise Ratio Acc Recall
CLIP (w/o DA) 0.2 93.88 97.49
CLIP (w DA) 0.2 97.68 97.18
CLIP (w/o DA) 0.5 93.91 99.35
CLIP (w DA) 0.5 98.14 99.24
@@
Model Time Extra Time
@@
4.6 Computational Cost Analysis
Cost in Pre-training.
To evaluate the practicality of OSA in a real-world pre-training scenario, we estimate the additional computational cost for processing 1 billion data points. Using an NVIDIA RTX 3090 with an inference batch size of 4096, utilizing approximately 24 GB of GPU memory, processing the MS-COCO dataset consisting of 566,435 pairs takes approximately 153 seconds. At this inference rate, processing 1 billion data points would require approximately 75 hours on a single RTX 3090. This cost is negligible within the context of large-scale pre-training, especially when leveraging multiple GPUs for parallel inference.
Time Cost Comparison.
To further examine the computational efficiency of OSA compared to other anti-noise techniques, we evaluate training time against two representative approaches: CLIP and NPC. CLIP, which serves as the baseline, is trained directly without any additional technique. NPC, the current state-of-the-art, also uses CLIP as its backbone but applies an anti-noise technique by estimating the negative impact of each sample, necessitating double backward passes. The training time comparison, presented in Table. 11, shows that our method introduces only a minimal increase in training time compared to direct training, requiring just one-tenth of the additional time needed by NPC. This highlights the efficiency of OSA, making it well-suited for large-scale robust training.
5 Conclusion
Broader Impacts.
In this work, we investigated the possibility of anti-noise in practical large-scale training. We introduced a novel model-agnostic anti-noise paradigm with advantages such as task transferability, model adaptability, and low computational overhead. By leveraging the properties of high-dimensional spaces, we found a robust and effective boundary for distinguishing between noisy and clean samples. Through rigorous theoretical analysis and comprehensive experimentation, we validated the efficacy and robustness of OSA for general noise mitigation. Although our primary objective is to adapt to practical large-scale training, OSA also achieves SOTA performance in standard anti-noise settings. To the best of our knowledge, this is the first work to explore anti-noise in practical large-scale training scenarios, as well as the first to propose a general anti-noise approach.
Limitations and Future Works.
Appendix
Appendix A Details of Implementation and Datasets
Dataset Details.
MSCOCO is widely used for noisy cross-modal matching, with each image accompanied by five descriptive captions. Following the setting of [1], we utilize 113,287 images for training, 5,000 for validation, and 5,000 for testing. The Flickr30K dataset encompasses 31,783 image-text instances, each image paired with five textual annotations. Adhering to the NCR [1], we use 29,783 images for training and 1,000 images each for validation and testing. Regarding noise splits, following the NCR categorization, we conduct experiments at noise ratios of 0%, 20%, 40%, and 60%. CC120K is a real-world multimodal noisy dataset collected by [2] from the Internet, with about 3%-20% noise ratio. There are 118,851 image-text pairs for training, 1,000 for validation, and 1,000 for testing.
The Aircraft, Bird, and Car we used in the image classification task are three non-overlapping subsets of the WebFG-496 [3] dataset. WebFG-496 consists of 53,339 images, totaling 496 subcategories. This dataset is annotated using a webly supervised approach, which leverages resources from web search engines (e.g., Google Image Search Engine, Bing Image Search Engine) to expand the annotated image dataset.
For the image retrieval task, we conduct experiments on the CARS98N dataset under PRISM’s setting [5]. We utilize 9,558 car-related images sourced from text-based searches on Pinterest as the training set, and employ the remaining 98 categories from CARS, unsearched on Pinterest, as a clean test set. The dataset’s noise is inherently real-world, with its creators estimating a noise ratio of approximately 50%.
Implementation Details.
To demonstrate the effectiveness of the OSA, we incorporate an estimator, built around the core of CLIP, and re-weighting operations based on the Estimator’s outcomes into numerous downstream tasks. In the principal task of cross-modal image-text retrieval, we employ CLIP with ViT-B/32 as the baseline and target model by default. All experiments are conducted on a single RTX 3090 GPU using the AdamW optimizer. During both training phases, the model is trained for five epochs with a batch size of 256 and 500 warmup steps.
For the image retrieval task, we use CLIP ViT-B/32 as the baseline, with a batch size set to 128, an initial learning rate of 5e-6, and the number of epochs set to 10. Following the setup of the PRISM [5], we set the parameter for sampling positive examples by the random sampler of the dataloader to 4, and adjust the number of positive examples sampled per epoch to one-fourth of the original parameter according to the increase in batch size. In this task, we also adopt a two-stage training approach. The strictly clean in-domain training data for the first stage is obtained using a GMM model with a probability setting of 0.8.
Appendix B Related work
B.1 Noise Mitigation in Cross-Modal Matching
The cross-modal matching task [24, 25, 26, 27, 28] serves as a fundamental component in multimodal learning. However, the inherent difference in information density between these modalities leads to high annotation costs and inconsistent annotation quality, rendering cross-modal tasks particularly vulnerable to label noise. Some approaches explicitly identify and correct noisy samples through cross-prediction between concurrently trained dual models [1, 7, 29], while others [2, 9] implicitly estimate the probability of sample noise, reducing its training impact by adjusting the loss function. NCR [1] employs the memorization capacity of its counterpart model for simple clean samples to rectify the output results. BiCro [7] utilizes the consistency of similarity score distributions from a Siamese model ensemble on noisy data, alongside anchors modeled on the loss distribution via a Beta-Mixture-Model (BMM), to filter out noisy samples. NPC [2], deviating from the dual-model training schemes, introduces a two-stage single-model training approach that reduces training overhead by replacing two backward passes with one forward and one backward pass. Specifically, the first stage estimates the impact of potentially noisy samples on model performance by constructing a high-quality clean sample bank; the second stage then utilizes these estimates to reweight the loss function. However, current methods for distinguishing clean from noisy samples rely on numerous hyperparameters that are closely linked to dataset size and model capacity. This dependency not only limits their adaptability to various downstream tasks but also makes them challenging to deploy in real-world applications.
B.2 Noise Mitigation in Image Classification
Image classification is vulnerable to training data noise, due to varied noise types and strong model memorization. Noise in datasets manifests in two primary forms: synthetic alterations and those arising from real-world scenarios. The former typically involves shuffling the labels of a subset of the data or retaining the labels while introducing corresponding category images from external datasets. The latter entails substituting images for a random selection of data points with those sourced from image search engines. Existing approaches are categorized based on their operational focus: loss correction [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] and sample selection [41, 42, 43, 10, 44]. Loss correction methods typically incorporate a regularization term into the loss function, implicitly reweighting clean and noisy samples within the loss. Sample selection strategies, in contrast, explicitly differentiate between clean and noisy samples, applying distinct processing to each category during loss computation. Representative for the loss correction category, [31] aims to generalize ordinary Cross-Entropy loss and MAE loss by setting the loss threshold to iid and ood noisy samples. DivideMix [10] concurrently trains two networks, each utilizing the data partitioning from the other network to distinguish between clean and noisy samples based on loss values, thereby mitigating the influence of confirmation bias inherent within each network. PNP [41] framework employs a unified predictive network to estimate the in-distribution (iid), out-of-distribution (ood), and clean probabilities for a given sample. Co-training trained on a sample that has a lower loss, and with the different predictions by its siamese network.
B.3 Noise Mitigation in Image Retrieval.
Although image retrieval tasks focus on pairwise relationships, the noise predominantly originates from image categorization errors. Analogous to image classification tasks, this can be bifurcated into in-domain [45] and open-set noise [5]. In terms of task configuration, noise retrieval typically operates at the category level, treating images within the same category as positive instances. PRISM [5] tries to find noisy image samples by finding the outliers score in the whole similarity matrix from the same category. The generalization ability of the image feature is ensured by a broader query bank restored multi-view of it. TITAN [46] utilizes prototypes to be representative of the anchor of the clean and noisy samples and then generates synthetic samples by a combination of prototypes for substitution of noisy samples. T-SINT [47] utilizes more negative samples by the interaction between noisy samples and negative samples that belong to another category.
Appendix C Proofs
C.1 Proof of High-dimensional Orthogonality
$\varsigma=\frac{1}{\sqrt{1024}}=\frac{1}{32},$ (8)
and
$P(-0.1\leq\cos(u,v)\leq 0.1)=\Phi\left(\frac{0.1}{1/32}\right)-\Phi\left(\frac{-0.1}{1/32}\right)\approx 0.9986.$ (9)
C.2 Proof of Theorem 1
In the Section. 2.2, we propose that Theorem 1 about the relative relationship of pairs in the original entire space, will not change after transmitting to the narrow cone space of the trained model, and there is always a boundary $r$ concentrated on most random vectors.
To prove this Theorem, we first introduce a useful lemma of monotonicity of cosine similarity proposed by [14], indicating that the cosine similarity between two vectors increases with a high probability after one feedforward computation consisting of a linear transformation and ReLU computation.
Lemma 1.
Proof of Theorem. 1.
Let $\mathbbm{R}^{d_{in}}$ be the original space before being transmitted in a neural network. Suppose $u,v\in\mathbbm{R}^{d_{in}}$ are any two random vectors with $\cos(u,v)\approx 0$. $u_{c},v_{c}\in\mathbbm{R}^{d_{in}}$ is a pair of clean vectors with $\cos(u_{c},v_{c})>0$, while $u_{n},v_{n}\in\mathbbm{R}^{d_{in}}$ is a noisy pair with $\cos(u_{n},v_{n})<0$. Given a Neural Network $F(x)={f_{t}(f_{t-1}(\dots f_{2}(f_{1}(x))))}\in\mathbbm{R}^{d_{out}}$ with $t$ layers. $f_{i}(x)=\sigma_{i}(\mathbf{W}_{i}x+\mathbf{b}_{i})$ denotes $i^{th}$ layer, where $\sigma(\cdot)$ indicates activation function. $\mathbf{W}_{i}\in\mathbbm{R}^{d^{i}_{out}\times d^{i}_{in}}$ is a random weight matrix where each element $\mathbf{W}_{i}^{k,l}\sim\mathcal{N}(0,1/d_{out}^{i})$ for $k\in\left[d^{i}_{out}\right]$, $l\in\left[d^{i}_{in}\right]$, and $\mathbf{b}_{i}\in\mathbbm{R}^{d^{i}_{out}}$ is a random bias vector such that $\mathbf{b}^{k}_{i}\sim\mathcal{N}(0,1/d^{i}_{out})$ for $k\in\left[d^{i}_{out}\right]$. We would like to prove that there are always be a boundary $\beta$, satisfying:
$\cos(F(u_{n}),F(v_{n}))<\cos(F(u),F(v))\approx\beta<\cos(F(u_{c}),F(v_{c})),$ (11)
which is equivalent to proving.
$\cos(f_{i}(u_{n}),f_{i}(v_{n}))<\cos(f_{i}(u),f_{i}(v))\approx\beta_{i}<\cos(f_{i}(u_{c}),f_{i}(v_{c})),$ (12)
where $\beta_{i}$ is the boundary of $i^{th}$ layer.
We first consider the cosine similarity between $u$ and $v$ as:
$\cos(u,v)=\frac{u\cdot v}{\|u\|\|v\|}.$ (13)
$\cos(\mathbf{W}_{i}u+\mathbf{b}_{i},\mathbf{W}_{i}v+\mathbf{b}_{i})=\frac{(\mathbf{W}_{i}u+\mathbf{b}_{i})\cdot(\mathbf{W}_{i}v+\mathbf{b}_{i})}{\|\mathbf{W}_{i}u+\mathbf{b}_{i}\|\|\mathbf{W}_{i}v+\mathbf{b}_{i}\|}.$ (14)
$\mathbb{E}\left[\mathbf{b}_{i}\right]=0,$ (15)
$\mathbb{E}\left[(\mathbf{W}_{i}u+\mathbf{b}_{i})\cdot(\mathbf{W}_{i}v+\mathbf{b}_{i})\right]=\mathbb{E}\left[(\mathbf{W}_{i}u\cdot\mathbf{W}_{i}v)\right]=\sum^{n}_{i=1}\sum^{n}_{i=1}\frac{1}{d^{i}_{out}}u_{k}v_{k}=\frac{1}{d^{i}_{out}}(u\cdot v).$ (16)
Additionally, we have
and
$\cos(\mathbf{W}_{i}u+\mathbf{b}_{i},\mathbf{W}_{i}v+\mathbf{b}_{i})$ $\approx\frac{\mathbb{E}[\mathbf{W}_{i}u\cdot\mathbf{W}_{i}v]}{\sqrt{\mathbb{E}[\|\mathbf{W}_{i}u+\mathbf{b}_{i}\|^{2}]\mathbb{E}[\|\mathbf{W}_{i}v+\mathbf{b}_{i}\|^{2}]}}$ (20)
$=\cos(u,v).$
$\cos(\mathbf{W}_{i}u_{n}+\mathbf{b}_{i},\mathbf{W}_{i}v_{n}\mathbf{b}_{i})<\cos(\mathbf{W}_{i}u+\mathbf{b}_{i},\mathbf{W}_{i}v+\mathbf{b}_{i})<\cos(\mathbf{W}_{i}u_{c}+\mathbf{b}_{i},\mathbf{W}_{i}v_{c}+\mathbf{b}_{i}).$ (21)
$\cos(f_{i}(u_{n}),f_{i}(v_{n}))<\cos(f_{i}(u),f_{i}(v))<\cos(f_{i}(u_{c}),f_{i}(v_{c})).$ (22)
$\cos(f_{i}(u_{n}),f_{i}(v_{n}))<\cos(f_{i}(u),f_{i}(v))\approx\beta_{i}<\cos(f_{i}(u_{c}),f_{i}(v_{c})).$ (23)
After transmitting each layer, Eq. 23 are always satisfied. When transmitting a neural network with $t$ layers, we have
$\cos(F(u_{n}),F(v_{n}))<\cos(F(u),F(v))\approx\beta<\cos(F(u_{c}),F(v_{c})).$ (24)
(a) parameters
(b) 128-th dim of image
(c) 256-th dim of image
(d) 512-th dim of image
(e) 128-th dim of text
(f) 256-th dim of text
(g) 512-th dim of text
C.3 Proof of Orthogonality Validity in Cone Space
To explore this, we first analyze the distribution of several dimensions of image and text features from the CC120K dataset, as illustrated in Figure. 3. The results reveal that all vector dimensions, including trained parameters, exhibit a Gaussian distribution with near-zero means. If the dimensions of the trained embedding space follow Gaussian distributions, the process of selecting random vectors within this space would be analogous to that of the original space, thereby preserving the orthogonality property. Here, we present the following theorem: The output features of large-scale models tend to Gaussian distribution. The detailed theorem and proof are provided below.
Theorem 2 (Output features tends to Gaussian).
Proof of Theorem. 2.
$\mathbb{E}\left[\sum_{j=1}^{d_{in}^{k}}w_{ij}^{k}x_{j}^{k}\right]=\sum_{j=1}^{d_{in}^{k}}\mathbb{E}[w_{ij}^{k}]\mathbb{E}[x_{j}^{k}]=\sum_{j=1}^{d_{in}^{k}}(0\times\mathbb{E}[x_{j}^{k}])=0.$ (25)
For variance, since $w_{ij}^{k}$ and $x_{j}^{k}$ are independent, we have:
$\text{Var}\left(\sum_{j=1}^{d_{in}^{k}}w_{ij}^{k}x_{j}^{k}\right)$ $=\sum_{j=1}^{d_{in}^{k}}\text{Var}(w_{ij}^{k}x_{j}^{k})=\sum_{j=1}^{d_{in}^{k}}\mathbb{E}\left[(w_{ij}^{k})^{2}(x_{j}^{k})^{2}\right]$ (26)
$=\sum_{j=1}^{d_{in}^{k}}\mathbb{E}\left[(w_{ij}^{k})^{2}\right]\mathbb{E}\left[(x_{j}^{k})^{2}\right]$
$=\sum_{j=1}^{d_{in}^{k}}\sigma_{w^{k}}^{2}\left(\text{Var}(x_{j}^{k})+(\mathbb{E}[x_{j}^{k}])^{2}\right)$
Since $w_{ij}^{k}$ are independently distributed Gaussian random variables, and $x_{j}^{k}$ has a known mean and variance, the sum of $w_{ij}^{k}x_{j}^{k}$ can apply to a generalized Central Limit Theorem. We have
[…]
which is equivalent to
[…]
Therefore,
[…]
Due to $b^{k}\sim\mathcal{N}(0,\sigma^{2}_{b})$, we finally get
[…]
Although activation functions truncate the Gaussian distribution after each linear layer, the samples still gradually approach a Gaussian distribution from the initial unknown distribution as they pass through the layers. Furthermore, because there is a fully connected layer ( layer) without an activation function before mapping to the final common space, the final feature distribution will approximate a Gaussian distribution, as follows:
Appendix D SDM Visualization

### v2

@@ D.1 Full Quantitative Results on Different Downstream Tasks
Table 7: Comparison on real-world noisy dataset CC120K and CC3M.
Method CC120K CC3M
@@
NPC 71.1 92.0 96.2 73.0 90.5 94.8 Out Of Memory
CLIP 68.8 87.0 92.9 67.8 86.4 90.9 42.41 66.70 75.56 42.45 67.83 76.46
+OSA 73.1 92.2 95.7 73.9 91.2 94.7 43.34 67.48 75.79 43.46 68.33 76.58
D.2 Full Results on Different Scoring Function
In Section 3.2.2, we introduce a scoring function designed to handle overlaps effectively and re-weight samples based on their cosine similarity. This section explores several scoring functions, including the Linear, Cosine, and High-degree functions. The functions are presented individually as follows:
Linear Function.
[…]
Cosine Function.
[…]
High-degree Function.
[…]
The results are presented in the Table. 8 indicates that the High-Degree Function outperforms the others across all evaluation metrics. This superior performance can be attributed to the rapid gradient changes near the decision boundary, which align better with the tendency of orthogonal boundaries. As a result, we adopt the High-Degree Function as our scoring function.
Table 8: Comparison of different scoring functions.
Noise ratio Method MS-COCO 1K MS-COCO 5K
i2t t2i i2t t2i
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
50% Linear 80.4 96.2 98.6 67.8 91.6 96.4 64.0 85.5 91.9 47.9 74.6 83.8
Cosine 80.8 96.3 98.5 67.7 91.6 96.3 64.4 86.2 92.3 48.0 74.9 83.9
High-Degree 81.4 96.5 98.6 68.4 92.0 96.6 64.7 86.8 92.4 48.6 75.9 84.6
D.3 Full Results on OSA Transferability
Results on transferring OSA to other noise mitigation methods. We apply OSA to give guidance to other noise mitigation methods to prove that our framework follows a plug-and-play style and can be easily fit into existing methods. From Table. 9, we observe that even integrated with the previous SOTA method, OSA can still give effective guidance to it, thereby the generalization ability of our method can be further proved.
@@ D.3 Full Results on OSA Transferability
Noise Ratio Method i2t t2i
R@1 R@5 R@10 R@1 R@5 R@10
@@
Table 10: The results of different estimator types on noisy MS-COCO.
@@
CLIP (w DA) 81.4 96.5 98.6 68.4 92.0 96.6 64.7 86.8 92.4 48.6 75.9 84.6
Results for overall noise rank reliability. Some anti-noise methods, like NPC, also employ loss re-weighting for optimization. To assess whether our method assigns relatively smaller weights to noise than these methods, we first analyze the weights generated by NPC and OSA. Due to differences in weight scales across methods, a direct comparison is unfair. We therefore adopt a ranking-based approach, sorting weights in descending order and calculating the Mean Noise Rank to unify the scale. This metric evaluates whether smaller weights are consistently assigned to noisy samples relative to clean ones. Our experiments use 2,000 randomly selected samples from the MSCOCO dataset under two noise conditions: 20% noise (370 noisy samples) and 50% noise (953 noisy samples). The theoretical optimal Mean Noise Ranks, where all noisy weights are ranked last, are 1815.5 and 1524.0, respectively. Results presented in Table. 11 show that OSA achieves a higher Mean Noise Rank compared to NPC, demonstrating greater accuracy in re-weighting. Moreover, OSA’s rankings are nearly optimal (20% noise: 1809.1 for OSA vs. 1815.5 optimal; 50% noise: 1520.7 for OSA vs. 1524.0 optimal). This near-perfect alignment indicates that OSA effectively places almost all noisy samples behind the clean ones.
@@
Model Time Extra Cost
@@
Comparisons of extra cost. As mentioned in Sec. 1, existing noise mitigation methods tend to have extra steps before training the target model. Table. 12 shows the comparisons between OSA and NPC on the extra time cost. As NPC needs extra time to coarsely filter some relative clean samples and use them to warm up the model, OSA just needs a forward step to calculate the noise probability, which leads to less extra time cost.
Appendix E SDM Visualization